OpenStudy (anonymous):

What sine function represents an amplitude of 1, a period of 2π, a horizontal shift of π, and a vertical shift of −4?

OpenStudy (anonymous):

f(x) = sin (x − π) − 4 f(x) = sin (πx) + 4 f(x) = sin (x + pi over 2) + 4 f(x) = sin (pi over 2x) − 4